OpenStudy (anonymous):

I really need some help with this problem, I have an exam and I cannot solve for the points. Find the point on the line -6 x + 5 y - 6 =0 which is closest to the point ( 4, -3 ).

OpenStudy (anonymous):

i did this for someone earlier on

OpenStudy (anonymous):

Would you please help me with it, I've been trying to solve it for a long time but I keep getting stuck and every answer I get is wrong, its very frustrating

OpenStudy (anonymous):

the slope of this equation is 6/5 and the line perpendicular to it has slope -5/6 and it must pass thru the point (4, -3) so its equation is y -(-3) = (-5/6)(x - 4) y+3 = (-5/6)(x - 4) multiply thru by 6 6y + 18 = -5x + 20 6y + 5x = 2.......(a) other equation is -6 x + 5 y - 6 =0 rearrange 5y - 6x = 6.........(b) WHERE THE lines of equations (a) and (b) intersect will be the shortest distance required we now need to solve system of equations (a) and (b) multiply (a) by 5 and (b) by 6: 30y + 25x = 10 30y - 36x = 36 subtract 61x = -26 x = -26/61 plug x into equation (a) 6y + 5* (-26/61) = 2 y = 2+ 130/61 = 252/61 The point on the line whch is the shoetest ditance is therefore ( -26/61, 252/61)

OpenStudy (anonymous):

the fractions are pretty messy - i might have made a mistake in solving ! but i'm sure the method is right

OpenStudy (anonymous):

I just double checked and x is correct but y is 42/61. Thank you so so so so much for your help, i've been going nuts trying to figure out the answer and I didn't know what i was doing wrong

OpenStudy (anonymous):

ok - yes - you are right y = 42/61 - i just checked it out on wolframalpha

OpenStudy (anonymous):

yep - i forgot to divide 252/61 by 6 on the last line - this gives 42/61

OpenStudy (anonymous):

no problem you have helped me so much and your explanation was great, i wrote it all down and understood exactly how you got your answer. Thanx a million!

OpenStudy (anonymous):

yw